# Text4Seg: Reimagining Image Segmentation as Text Generation

### v1

@@ Appendix D Additional Qualitative Results
Fig. 17 presents additional examples of Text4Seg performing open-vocabulary segmentation. Notably, Text4Seg demonstrates its ability to segment not only common large objects but also small objects effectively, such as the person and boat on the river. This versatility highlights Text4Seg’s proficiency in accurately identifying and segmenting a wide range of object sizes. Fig. 18 illustrates the multi-object segmentation capabilities of Text4Seg. It is evident that Text4Seg successfully segments all identified objects within the image, showcasing its strong ability to handle multiple objects in complex scenarios. This performance highlights its robustness and effectiveness in accurately distinguishing various elements within a single scene.
@@
Fig. 19 presents an example where Text4Seg is used for image captioning, single-object segmentation, and multi-object segmentation. Additionally, Fig. 20 compares the image reasoning capabilities of Text4Seg with the original LLaVA-1.5. While maintaining similar reasoning abilities, our proposed Text4Seg extends functionality by enabling segmentation tasks.

### v2

@@ Appendix E Additional Qualitative Results
Figure 17 presents additional examples of Text4Seg performing open-vocabulary segmentation. Notably, Text4Seg demonstrates its ability to segment not only common large objects but also small objects effectively, such as the person and boat on the river. This versatility highlights Text4Seg’s proficiency in accurately identifying and segmenting a wide range of object sizes. Figure 18 illustrates the multi-object segmentation capabilities of Text4Seg. It is evident that Text4Seg successfully segments all identified objects within the image, showcasing its strong ability to handle multiple objects in complex scenarios. This performance highlights its robustness and effectiveness in accurately distinguishing various elements within a single scene.
@@
Figure 19 presents an example where Text4Seg is used for image captioning, single-object segmentation, and multi-object segmentation. Additionally, Fig. 20 compares the image reasoning capabilities of Text4Seg with the original LLaVA-1.5. While maintaining similar reasoning abilities, our proposed Text4Seg extends functionality by enabling segmentation tasks.